# Plenoptic Cameras; Light Field; Calibration; Refocusing; Simulation

### 20 Feb 2024 v1

@@ 6 Conclusion and Limitations
Further limitations are concerned with the formal lens model used for the error deduction. First, the microlenses in our model are still formally regarded as thin lenses. While Hahne et al. Hahne et al. (2016) use a thick lens model with explicit microlens principal planes, it was decided to leave this aspect out of the theoretical discussion for reasons of clarity and comprehensibility. However, the microlens thickness was factored in while performing the experiments and appendix D shows, that an extended model does not change the equations deduced in section 2 and section 3.
@@ Appendix B Revisited Literature - Notation Transfer
In the following sections, the notation of Dansereau et al. (2013) and Pertuz et al. (2018) are transferred into the notation used in this work.
B.1 Dansereau et al.Dansereau et al. (2013)
@@
Appendix C Evaluation Setups
In the following tables 6 and 7, the specific lens models, associated patents and the SPC configurations are listed. Note, that $H_{\text{cam}}$ and the exit pupil location are measured from the main lens aperture position and have a positive sign for these planes being located on the sensor side of the aperture. On the other hand, $H_{\text{scene}}$ is measured from the aperture with a positive sign if located on the scene-side of the aperture. $X$ is measured from the position of $H_{\text{cam}}$ as shown in figure 2 and all remaining properties are assumed to be unsigned distance values. All mentioned properties were measured via a 2d ray tracer which we also integrated into the Blender add-on to facilitate fast automatic plenoptic camera reconfigurations.
@@ Appendix C Evaluation Setups
Table 6: Exact model names of the simulated main lenses and associated patents specifying the lens element parameters such as curvature, height and index of refraction. DE, JP and US specify the country of the patent application, i.e. Germany, Japan or USA.Claff
@@
Table 7: SPC properties for the evaluation setups.
Appendix D Thick Microlenses

### 05 Apr 2024 v2

@@ 6 Conclusion and Limitations
Further limitations are concerned with the formal lens model used for the error deduction. First, the microlenses in our model are still formally regarded as thin lenses. While Hahne et al. Hahne et al. (2016) use a thick lens model with explicit microlens principal planes, it was decided to leave this aspect out of the theoretical discussion for reasons of clarity and comprehensibility. However, the microlens thickness was factored in while performing the experiments and appendix C shows, that an extended model does not change the equations deduced in section 2 and section 3.
@@ Appendix B Revisited Literature - Notation Transfer
In the following sections, the notation of Hahne et al. (2016), Dansereau et al. (2013) and Pertuz et al. (2018) are transferred into the notation used in this work.
B.1 Hahne et al.Hahne et al. (2016)
Description Hahne et al. Hahne et al. (2016) Our
Main lens focal length $f_{U}$ $f_{M}$
Distance $H_{\text{cam}}$ to MLA $b_{U}$ $d$
Distance $H_{\text{cam}}$ to virtual image $b_{U^{\prime}}$ i
Distance $H_{\text{scene}}$ to scene object $a_{U^{\prime}}$ o
Distance MLA to exit pupil $d_{A^{\prime}}$ $d-X$
Distance MLA to virtual image $d^{\prime}_{a}$ $d-\textbf{i}$
ML focal length $f_{s}$ $f_{m}$
ML diameter $\Delta s$ $d_{\text{ML}}$
Pixel size $\Delta u$ $s_{px}$
Table 3: General notation in Hahne et al. Hahne et al. (2016) and our work.
In addition to the general camera properties listed in table 3, Hahne et al. Hahne et al. (2016) introduce the notation in table 4 to exemplarily describe the camera-side rays for a horizontal cross section of the plenoptic camera model.
Description Hahne et al. Hahne et al. (2016)
Horizontal number of microlenses $J$
Microlens index $j$
Microlens center of lens $j$ $s_{j}=\left(j-\frac{J-1}{2}\right)\cdot\Delta s$
Microlens image center of lens $j$ $u_{c,j}=\frac{s_{j}}{d_{A^{\prime}}}\cdot f_{s}+s_{j}$
Position of i-th neighbor pixel of MIC $u_{c+i,j}=u_{c,j}+i\cdot\Delta u$
Slope of ray from i-th neighbor through ML center $m_{c+i,j}=\frac{s_{j}-u_{c+i,j}}{f_{s}}$
Table 4: Further notation in Hahne et al. Hahne et al. (2016) without direct equivalents in our setup.
With these the linear ray function originating in a pixel at the sensor position $u_{c+i,j}$ and running through the corresponding microlens center $s_{j}$ can be described via
$f_{c+i,j}(z)$ $=m_{c+i,j}\cdot z+s_{j}=\frac{s_{j}-u_{c+i,j}}{f_{s}}\cdot z+s_{j}$
$=\frac{s_{j}-(u_{c,j}+i\cdot\Delta u)}{f_{s}}\cdot z+s_{j}=\frac{s_{j}-\left(\frac{s_{j}}{d_{A^{\prime}}}\cdot f_{s}+s_{j}+i\cdot\Delta u\right)}{f_{s}}\cdot z+s_{j}.$ (34)
By assuming an MLA with a microlens center located at the main lens optical axis, the ray originating from the center of the central microlens image going through the microlens center $s_{0}=0$ can be described by
$f(z):=f_{c,0}(z)=\frac{s_{0}-\left(\frac{s_{0}}{d-X}\cdot f_{m}+s_{0}+0\cdot s_{px}\right)}{f_{m}}\cdot z+s_{0}=0.$ (35)
Given a second pixel position at distance $d_{\text{MLI}}+\hat{S}$ from the central pixel, this pixel’s index in the microlens image $j=1$ is given by $i=\frac{\hat{S}}{s_{px}}$. Accordingly, the ray originating from this pixel running through the microlens center $s_{1}=d_{\text{ML}}$ is given by
$\tilde{f}(z):=f_{c+i,1}(z)$ $=\frac{s_{1}-\left(\frac{s_{1}}{d-X}\cdot f_{m}+s_{1}+\frac{\hat{S}}{s_{px}}\cdot s_{px}\right)}{f_{m}}\cdot z+s_{1}$
$=\frac{d_{\text{ML}}-\left(\frac{d_{\text{ML}}}{d-X}\cdot f_{m}+d_{\text{ML}}+\hat{S}\right)}{f_{m}}\cdot z+d_{\text{ML}}$
$=\frac{d_{\text{ML}}-\left(d_{\text{MLI}}+\hat{S}\right)}{f_{m}}\cdot z+d_{\text{ML}}.$ (36)
B.2 Dansereau et al.Dansereau et al. (2013)
@@
Appendix C Thick Microlenses
Appendix D Evaluation Setups
In the following tables 8, 9 and 10, the specific lens models, associated patents and the SPC configurations are listed.
@@ Appendix D Evaluation Setups
Table 8: Model names of the simulated main lenses and associated patents specifying the lens element parameters. DE, JP and US specify the country of the patent application.Claff
@@
Table 9: SPC properties for the evaluation setups with finite focus distance.
@@
Table 10: SPC properties for the evaluation setups with infinite focus distance.
Note, that $H_{\text{cam}}$ and the exit pupil location are measured from the main lens aperture position and have a positive sign for these planes being located on the sensor side of the aperture. On the other hand, $H_{\text{scene}}$ is measured from the aperture with a positive sign if located on the scene-side of the aperture. $X$ is measured from the position of $H_{\text{cam}}$ as shown in figure 2 and all remaining properties are assumed to be unsigned distance values. All mentioned properties were measured via a 2d ray tracer which we also integrated into the Blender add-on to facilitate fast automatic plenoptic camera reconfigurations.